### v1

# Enhancing Diffusion Model Stability for Image Restoration via Gradient Management

Abstract.
Diffusion models have shown remarkable promise for image restoration by leveraging powerful priors. Prominent methods typically frame the restoration problem within a Bayesian inference framework, which iteratively combines a denoising step with a likelihood guidance step. However, the interactions between these two components in the generation process remain underexplored. In this paper, we analyze the underlying gradient dynamics of these components and identify significant instabilities. Specifically, we demonstrate conflicts between the prior and likelihood gradient directions, alongside temporal fluctuations in the likelihood gradient itself. We show that these instabilities disrupt the generative process and compromise restoration performance. To address these issues, we propose Stabilized Progressive Gradient Diffusion (SPGD), a novel gradient management technique. SPGD integrates two synergistic components: (1) a progressive likelihood warm-up strategy to mitigate gradient conflicts; and (2) adaptive directional momentum (ADM) smoothing to reduce fluctuations in the likelihood gradient. Extensive experiments across diverse restoration tasks demonstrate that SPGD significantly enhances generation stability, leading to state-of-the-art performance in quantitative metrics and visually superior results. Code is available at here.
copyright: acmlicensedjournalyear: 2025conference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelanddoi: XXXXXXX.XXXXXXXsubmissionid: 7272ccs: Computing methodologies Computer vision problemsccs: Computing methodologies Machine learning
Refer to captionInputDPSOursReference
Figure 1. Illustrative comparison of diffusion-based restoration results: DPS (without gradient management), Ours (with the proposed gradient management), and the reference. Our method demonstrates improved restoration quality, particularly in challenging restoration scenarios.
1. Introduction
Image restoration (Zhang et al., 2017; Lehtinen et al., 2018; Liang et al., 2021), encompassing tasks such as inpainting (Yu et al., 2018; Elharrouss et al., 2020), deblurring (Chen et al., 2024; Wu et al., 2024c), and super-resolution (Saharia et al., 2022b; Wang et al., 2024b), is a fundamental problem in computer vision and image processing. The goal is to recover a high-quality, clean image $\mathbf{x}_{0}$ from a degraded measurement $\mathbf{y}$, which is often modeled by the following equation:
(1) $\mathbf{y}=\mathcal{A}(\mathbf{x}_{0})+\mathbf{n},$
where $\mathcal{A}$ represents a known degradation operator (e.g., a blurring kernel, a downsampling matrix, or a masking operator), and $\mathbf{n}\sim\mathcal{N}(\mathbf{0},{\sigma_{y}^{2}}\mathbf{I})$ represents additive noise. Traditional approaches to solve this ill-posed inverse problem (Milanfar, 2012; Zhang et al., 2017; Tai et al., 2017; Zhang et al., 2021) often rely on handcrafted priors and regularization terms, which may not fully capture the complex statistics of natural images.
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b) have emerged as a transformative force in generative modeling, achieving state-of-the-art results in image synthesis (Rombach et al., 2022; Poole et al., 2022; Saharia et al., 2022a; Meng et al., 2023). Their remarkable ability to model complex data distribution and generating realistic images, has naturally led to significant interest in adapting them for image restoration (Choi et al., 2021; Wen et al., 2024; Lugmayr et al., 2022; Carrillo et al., 2023; Cao et al., 2024). Among these methods, a dominant paradigm leverages the Bayesian framework (Song et al., 2022; Chung et al., 2023; Song et al., 2023; Dou and Song, 2024), guiding the reverse diffusion process using the conditional score $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{y})$. This score elegantly decomposes via Bayes’ theorem into two key components:
(2) $\underbrace{\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{y})}_{\text{Conditional Score}}=\underbrace{\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})}_{\text{Prior Score}}+\underbrace{\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}|\mathbf{x}_{t})}_{\text{Likelihood Guidance}}.$
Here, the prior score enforces conformity to the learned distribution of natural images, while the likelihood guidance ensures consistency with the observed measurement $\mathbf{y}$.
Existing methods (Chung et al., 2023; Song et al., 2023; Peng et al., 2024; Wu et al., 2024a) typically apply the gradients corresponding to the prior score and likelihood guidance iteratively, without explicitly accounting for their potentially detrimental interactions. However, despite the success of these approaches, we argue that their performance is limited by inherent conflicts and instabilities within the gradient dynamics governing the reverse process. Our detailed analysis in Section 3, which examines the reverse update from a gradient decomposition perspective, reveals two critical issues:
  1. (1)
    Gradient Conflict: The gradient enforcing data consistency (likelihood guidance) often conflicts with the gradient enforcing the natural image prior in direction, particularly in the early stages of generation.
  2. (2)
    Gradient Fluctuation: The likelihood guidance gradient exhibits significant temporal instability, fluctuating abruptly between consecutive timesteps.
Such issues can hinder the effective combination of prior knowledge and data fidelity, leading to suboptimal updates and ultimately compromising generation quality. Addressing these instabilities is essential for unlocking the full potential of diffusion models in demanding restoration tasks.
This paper introduces Stabilized Progressive Gradient Diffusion (SPGD), a novel gradient management technique designed to address these identified instabilities. SPGD integrates two synergistic components into the reverse diffusion step:
  1. (1)
    Progressive Likelihood Warm-Up: To mitigate gradient conflict, we introduce multiple small, iterative updates guided only by the likelihood gradient before each main denoising step. This allows the estimate to progressively adapt to measurement constraints, reducing potential directional clashes with the prior gradient.
  2. (2)
    Adaptive Directional Momentum (ADM) Smoothing: To counteract likelihood gradient instability, we employ ADM smoothing. This technique dynamically adjusts momentum based on directional consistency, effectively dampening erratic fluctuations and promoting more stable update guidance from the likelihood term.
These components work in concert to stabilize the reverse trajectory, leading to more effective and robust optimization, even with fewer computational resources. Extensive experiments across diverse restoration tasks demonstrate that SPGD significantly enhances reconstruction quality, yielding substantial improvements in restoration metrics alongside visually superior results.
In summary, our primary contributions are:
  • We present a novel gradient-centric perspective on the reverse process in diffusion models for image restoration, identifying the critical, yet previously under-addressed, issues of gradient conflict and fluctuation.
  • We propose Stabilized Progressive Gradient Diffusion (SPGD), a novel gradient management technique incorporating a progressive warm-up phase and momentum-based smoothing to mitigate these issues.
  • We conduct extensive experiments on various image restoration tasks across two different datasets, demonstrating that SPGD significantly improves reconstruction quality both quantitatively and qualitatively.
2. Background
(a) Gradient angles dynamics without smoothing.
(b) Gradient angles dynamics with ADM smoothing.
Figure 2. Angular relationships of gradients during the reverse process across different tasks, illustrating the gradient angles dynamics (a) without smoothing and (b) with our ADM smoothing. Blue lines: angle between $\mathbf{g}_{l}$ and $\mathbf{g}_{d}$. Orange lines: angle between $\mathbf{g}_{l}(\mathbf{x}_{t})$ and $\mathbf{g}_{l}(\mathbf{x}_{t-1})$. Green lines: angle between $\mathbf{g}_{d}(\mathbf{x}_{t})$ and $\mathbf{g}_{d}(\mathbf{x}_{t-1})$.
2.1. Diffusion Models
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b), also known as score-based models, operate by defining a stochastic process that gradually perturbs data samples into a noise distribution and subsequently learns to reverse this process to generate samples from the desired data distribution. Consider a $T$-step forward process, where the noised sample $\mathbf{x}_{t}\in\mathbb{R}^{d}$ at time step $t$ can be modeled from the previous state $\mathbf{x}_{t-1}$:
(3) $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}),$
where $\mathcal{N}$ denotes the Gaussian distribution, and $\beta_{t}$ is a pre-defined parameter increasing with $t$. Through reparameterization, $\mathbf{x}_{t}$ can be directly expressed in terms of the original data $\mathbf{x}_{0}$ via:
(4) $q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I}),$
where $\bar{\alpha}_{t}=\prod_{i=0}^{t}\alpha_{i}$ and $\alpha_{t}=1-\beta_{t}$.
The training objective of diffusion models is typically defined as:
(5) $\mathbf{L}({\mathbf{\theta}})=\mathbb{E}_{\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t\sim\mathbb{U}(\{1,...,T\})}[\|\bm{\epsilon}-\bm{\epsilon}_{\mathbf{\theta}}(\mathbf{x}_{t},t)\|^{2}],$
where $\mathbf{x}_{0}$ is sampled from training data and $\mathbf{x}_{t}\sim q(\mathbf{x}_{t}|\mathbf{x}_{0})$. Once the network $\bm{\epsilon}_{\mathbf{\theta}}(\mathbf{x}_{t},t)$ is trained, a clean sample can be derived by evaluating the generative reverse process step by step.
It is worth noting that one can adopt DDIM (Song et al., 2021a) to perform the reverse sampling, as it preserves the same marginal distributions $q(\mathbf{x}_{t}|\mathbf{x}_{0})$ as DDPM. The DDIM sampling procedure can be formulated as:
(6) $\mathbf{x}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\mathbf{\bm{\epsilon}}_{\mathbf{\theta}}(\mathbf{x}_{t},t)+\sigma_{t}\bm{\epsilon},$
where $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})=(\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{\mathbf{\theta}}(\mathbf{x}_{t},t))/{\sqrt{\bar{\alpha}_{t}}}$ is the predicted clean image given $\mathbf{x}_{t}$. The parameter $\sigma_{t}$ governs the variance of the reverse sampling process. Specifically, setting $\sigma_{t}=0$ yields a deterministic DDIM sampling, while $\sigma_{t}=\sqrt{\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}}$ recovers the original DDPM.
2.2. Diffusion-Based Solvers for Image Restoration
The strong generative power of diffusion models makes them excellent candidates for solving image restoration and inverse problem (Xia et al., 2023; Ye et al., 2024a; Zhang et al., 2024). To implement this, the reverse process is guided by the conditional score $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{y})$. As shown in Equation 2, this involves the unconditional score $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ and the likelihood term $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}|\mathbf{x}_{t})$.
The unconditional score $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ is directly related to $\frac{-\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)}{\sqrt{1-\bar{\alpha}_{t}}}$ (Song et al., 2021b). However, the likelihood term is intractable because $p_{t}(\mathbf{y}|\mathbf{x}_{t})$ is unknown, since there only exists an explicit connection between $\mathbf{y}$ and $\mathbf{x}_{0}$, not $\mathbf{x}_{t}$. To deal with this, a family of methods avoids explicit likelihood computation through interleaving optimization (Song et al., 2024; Zhu et al., 2023; Li et al., 2024; Wu et al., 2024b) or projecting the estimate onto a measurement-consistent subspace (Kawar et al., 2022; Chung et al., 2022; Wang et al., 2023; Cardoso et al., 2024).
The another category of methods direct compute the likelihood under some mild assumptions (Rout et al., 2023, 2024). Of these, Chung et al. (2023) propose Diffusion Posterior Sampling (DPS) to approximate the likelihood using a Laplacian approximation: $p_{t}(\mathbf{y}|\mathbf{x}_{t})\simeq p_{t}(\mathbf{y}|\hat{\mathbf{x}}_{0}(\mathbf{x}_{t}))$, $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})$ is the predicted clean image of $\mathbf{x}_{t}$ via Tweedie’s formula (Stein, 1981; Efron, 2011). Then the likelihood gradient can be calculated by:
(7) $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}|\mathbf{x}_{t})\simeq-{\zeta}\nabla_{\mathbf{x}_{t}}\|\mathbf{y}-\mathcal{A}(\hat{\mathbf{x}}_{0}(\mathbf{x}_{t}))\|_{2}^{2},$
where $\zeta$ is a tunable step size that controls the strength of the guidance. The complete DPS algorithm with DDIM deterministic update is shown below:
(8) $\mathbf{x}_{t-1}^{\prime}$ $=\sqrt{\bar{\alpha}_{t-1}}\big{(}\frac{\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{\mathbf{\theta}}(\mathbf{x}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\big{)}+\sqrt{1-\bar{\alpha}_{t-1}}\bm{\epsilon}_{\mathbf{\theta}}(\mathbf{x}_{t},t),$
$\mathbf{x}_{t-1}$ $=\mathbf{x}_{t-1}^{\prime}-{\zeta}\nabla_{\mathbf{x}_{t}}\|\mathbf{y}-\mathcal{A}(\hat{\mathbf{x}}_{0}(\mathbf{x}_{t}))\|_{2}^{2}.$
By iteratively updating via likelihood matching combined with denoising, one can effectively recover high-quality estimates of $\mathbf{x}_{0}$ from degraded observations $\mathbf{y}$. However, as we will show, the direct application of these updates can be unstable.
3. A Gradient Perspective on Reverse Process
3.1. Reverse Update Decomposition
Despite the empirical success of Bayesian guidance frameworks, the underlying behavior and interaction between the prior score and likelihood guidance warrant closer examination. In this section, we adopt a gradient-centric perspective (Dinh et al., 2023) to analyze the reverse update process.
Consider the DPS likelihood guidance combined with the deterministic DDIM update ($\sigma_{t}=0$), we rearrange the update from $\mathbf{x}_{t}$ to $\mathbf{x}_{t-1}$ to explicitly highlight the contributing gradient terms. The insightful decomposition (see Appendix A for derivation) is:
(9) $\mathbf{x}_{t-1}=$ $\underbrace{\frac{1}{\sqrt{\alpha_{t}}}\mathbf{x}_{t}}_{\text{fixed scaling}}-\underbrace{\left(\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{{\alpha}_{t}}}-\sqrt{1-\bar{\alpha}_{t-1}}\right)\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)}_{\text{denoising gradient}~{}(\mathbf{g}_{d})}$
$-{\zeta}\underbrace{\nabla_{\mathbf{x}_{t}}\|\mathbf{y}-\mathcal{A}(\hat{\mathbf{x}}_{0}(\mathbf{x}_{t}))\|_{2}^{2}}_{\text{likelihood gradient}~{}(\mathbf{g}_{l})},$
Here, we explicitly define:
  • $\mathbf{g}_{d}(\mathbf{x}_{t})$: The denoising gradient, derived from the learned prior $\bm{\epsilon}_{\bm{\theta}}$, guiding the sample towards the natural image manifold.
  • $\mathbf{g}_{l}(\mathbf{x}_{t})$: The likelihood gradient, enforcing consistency with the measurement $\mathbf{y}$ via the reconstruction error.
The fixed scaling term primarily adjusts the magnitude of $\mathbf{x}_{t}$. The core dynamics influencing the image content arise from the interplay between $\mathbf{g}_{d}$ and $\mathbf{g}_{l}$. While previous works (Chung et al., 2023; Song et al., 2023; Peng et al., 2024; Wu et al., 2024a) implicitly combine these effects, analyzing their individual dynamics reveals potential sources of performance degradation.
3.2. Gradient Conflicts and Fluctuations
We empirically examine the angular relationships between $\mathbf{g}_{d}$ and $\mathbf{g}_{l}$, as well as the temporal consistency of each gradient. The results are shown in Figure 2 (see Appendix E for more results). Our analysis reveals two critical issues:
(1) Gradient Conflict: The angle between $\mathbf{g}_{d}(\mathbf{x}_{t})$ and $\mathbf{g}_{l}(\mathbf{x}_{t})$ (blue lines in Figure 2) frequently diverges from 90 degrees in the early stage, implying that the prior direction conflicts with the data constraint direction. While these conflicts diminish as $t$ decreases 111two randomly chosen vectors tend to be orthogonal in high dimensions (Ledoux, 2001; Talagrand, 1995), their presence in the early stages can disrupt the image structure formation (Singh et al., 2022; Everaert et al., 2024), a crucial phase for overall reconstruction quality.
(2) Likelihood Gradient Fluctuation: The temporal consistency between $\mathbf{g}_{l}(\mathbf{x}_{t})$ and $\mathbf{g}_{l}(\mathbf{x}_{t+1})$ (orange lines in Figure 2) often exhibits pronounced instability in the intermediate stages. We speculate that $\mathbf{g}_{l}$ is derived from $\|\mathbf{y}-\mathcal{A}(\hat{\mathbf{x}}_{0}(\mathbf{x}_{t}))\|_{2}^{2}$, which can be highly sensitive to small changes in the estimate $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})$. These erratic fluctuations hinder smooth convergence towards data consistency and can potentially amplify errors over time.
In contrast, the denoising gradient $\mathbf{g}_{d}$ derived from the well-trained $\bm{\epsilon}_{\bm{\theta}}$ demonstrates remarkable temporal stability (visualized by the green lines in Figure 2), providing a consistent guidance from the learned prior.
These observed conflicts and fluctuations in the gradient dynamics are not merely theoretical concerns; they manifest as tangible artifacts in reconstructed images, slower convergence rates, and quantifiable reductions in evaluation metrics (see Section 5). This motivates the development of a method that explicitly manages these gradient dynamics for more stable and effective restoration.
4. Proposed Method
Motivated by our analysis in Section 3, we propose Stabilized Progressive Gradient Diffusion (SPGD), a novel gradient management technique to enhance the stability and effectiveness of the reverse diffusion process. SPGD introduces two synergistic components: a progressive likelihood warm-up strategy and a momentum-based likelihood gradient smoothing technique.
4.1. Progressive Warm-Up Phase
To mitigate the conflicts between the denoising gradient $\mathbf{g}_{d}$ and the likelihood gradient $\mathbf{g}_{l}$, SPGD introduces a progressive warm-up phase for likelihood update before each primary denoising step.
Specifically, within a single conceptual reverse step from time $t$ to $t-1$, starting from the current estimate $\mathbf{x}_{t}^{(0)}=\mathbf{x}_{t}$, we perform $N$ iterative refinement steps focusing solely on the likelihood gradient using a reduced step size:
(10) $\mathbf{x}_{t}^{(j+1)}=\mathbf{x}_{t}^{(j)}-\frac{\zeta}{N}\widetilde{\mathbf{g}}_{l}(\mathbf{x}_{t}^{(j)}),\quad\text{for }j=0,1,\ldots,N-1,$
where $\widetilde{\mathbf{g}}_{l}(\mathbf{x}_{t}^{(j)})$ is the ADM-smoothed likelihood gradient (detailed in Equation 12), and $\zeta/N$ serves as a small learning rate for each warm-up step. This multi-step process allows the estimate $\mathbf{x}_{t}$ to gradually adapt towards satisfying the measurement constraint $\mathbf{y}$ before the potentially conflicting denoising gradient $\mathbf{g}_{d}$ is applied. Moreover, the small step size helps prevent drastic deviations from the learned data manifold (Alkhouri et al., 2024; Huang et al., 2022), fostering a more controlled adjustment.
After these $N$ warm-up iterations yield $\mathbf{x}_{t}^{(N)}$, the standard denoising update is applied to obtain the initial estimate for the next timestep, using $\mathbf{x}_{t}^{(N)}$ as input:
(11) $\mathbf{x}_{t-1}^{(0)}=\frac{1}{\sqrt{\alpha_{t}}}\mathbf{x}_{t}^{(N)}-\left(\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{{\alpha}_{t}}}-\sqrt{1-\bar{\alpha}_{t-1}}\right)\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t}^{(N)},t),$
which is a standard denoising update with diffusion model, i.e. DDIM in this paper. Given the observed stability of the denoising gradient $\mathbf{g}_{d}$, this multi-step warm-up mitigates the instability risk arising from potential conflicts, providing a more refined input $\mathbf{x}_{t}^{(N)}$ for the reliable denoising operation.
Algorithm 1 Stabilized Progressive Gradient Diffusion
0:  Measurements $\mathbf{y}$, degrade operator $\mathcal{A}(\cdot)$, network $\bm{\epsilon}_{\theta}(\cdot)$, parameters $\zeta$
1:  Initialize $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
2:  for $t=T,\dots,1$ do
3:     for $j=0,\dots,N-1$ do
4:        $\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})=\nabla_{\mathbf{x}_{t}^{(j)}}\|\mathbf{y}-\mathcal{A}(\hat{\mathbf{x}}_{0}(\mathbf{x}_{t}^{(j)}))\|_{2}^{2}$
5:        if j=0 then
6:           $\widetilde{\mathbf{g}}_{l}(\mathbf{x}_{t}^{(0)})=\mathbf{g}_{l}(\mathbf{x}_{t}^{(0)})$
7:        else
8:           $\alpha_{j}=(\text{sim}\big{(}\widetilde{\mathbf{g}}_{l}(\mathbf{x}_{t}^{(j-1)}),\,\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})\big{)}+1)/2,$
9:           $\widetilde{\mathbf{g}}_{l}(\mathbf{x}_{t}^{(j)})=\alpha_{j}\beta\,\widetilde{\mathbf{g}}_{l}(\mathbf{x}_{t}^{(j-1)})+(1-\alpha_{j}\beta)\,\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})$ // adaptive momentum-based gradient smoothing
10:        end if
11:        $\mathbf{x}_{t}^{(j+1)}\leftarrow\mathbf{x}_{t}^{(j)}-\frac{\zeta}{N}\widetilde{\mathbf{g}}_{l}(\mathbf{x}_{t}^{(j)})$ // warm-up likelihood update
12:     end for
13:     $\mathbf{g}_{d}(\mathbf{x}_{t}^{(N)})=\left({\sqrt{1-\bar{\alpha}_{t}}}/{\sqrt{{\alpha}_{t}}}-\sqrt{1-\bar{\alpha}_{t-1}}\right)\bm{\epsilon}_{\theta}(\mathbf{x}_{t}^{(N)},t)$
14:     $\mathbf{x}_{t-1}^{(0)}=\frac{1}{\sqrt{\alpha_{t}}}\mathbf{x}_{t}^{(N)}-\mathbf{g}_{d}(\mathbf{x}_{t}^{(N)})$ // DDIM denoising update
15:  end for
16:  return  $\mathbf{x}_{0}$
4.2. Momentum-Based Likelihood Gradient Smoothing
To counteract the temporal fluctuation observed in the likelihood gradient $\mathbf{g}_{l}$, SPGD incorporates an Adaptive Directional Momentum (ADM) smoothing technique. This produces the stabilized likelihood gradient $\widetilde{\mathbf{g}}_{l}$ by incorporating historical information, which is then used in the warm-up phase in Equation 10.
Standard momentum methods (Polyak, 1964; Sutskever et al., 2013; He et al., 2024; Wang et al., 2024a) use an exponentially weighted moving average to smooth gradients. However, constant momentum can slow adaptation when the optimal direction genuinely changes. Our proposed ADM enhances this by dynamically adjusting the momentum strength based on the directional consistency between the current gradient and the accumulated momentum.
Specifically, let $\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})=\nabla_{\mathbf{x}_{t}^{(j)}}\|\mathbf{y}-\mathcal{A}(\hat{\mathbf{x}}_{0}(\mathbf{x}_{t}^{(j)}))\|_{2}^{2}$ be the raw likelihood gradient at warm-up step $j$. The ADM-smoothed gradient $\widetilde{\mathbf{g}}_{l}(\mathbf{x}_{t}^{(j)})$ is computed recursively:
(12) $\widetilde{\mathbf{g}}_{l}(\mathbf{x}_{t}^{(j)})=\alpha_{j}\beta\,\widetilde{\mathbf{g}}_{l}(\mathbf{x}_{t}^{(j-1)})+(1-\alpha_{j}\beta)\,\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)}),$
with initialization $\widetilde{\mathbf{g}}_{l}(\mathbf{x}_{t}^{(0)})=\mathbf{g}_{l}(\mathbf{x}_{t}^{(0)})$. Here, $\beta\in[0,1)$ is the pre-defined base momentum coefficient. The key component is the adaptive weight $\alpha_{j}\in[0,1]$, defined as:
(13) $\alpha_{j}=\frac{\text{sim}\big{(}\widetilde{\mathbf{g}}_{l}(\mathbf{x}_{t}^{(j-1)}),\,\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})\big{)}+1}{2},$
where $\text{sim}(\mathbf{u},\mathbf{v})=\frac{\mathbf{u}\cdot\mathbf{v}}{\|\mathbf{u}\|\|\mathbf{v}\|}\in[-1,1]$, denotes the cosine similarity.
Theoretical intuition: When the current gradient $\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})$ aligns well ($\text{sim}(\cdot)\approx 1$) with the previous momentum $\widetilde{\mathbf{g}}_{l}(\mathbf{x}_{t}^{(j-1)})$, then $\alpha_{j}\approx 1$, and ADM behaves like standard momentum, smoothing the update by trusting the historical trend. Conversely, when the directions differ significantly ($\text{sim}(\cdot)\ll 1$), $\alpha_{j}$ decreases towards 0, effectively down-weighting the potentially outdated historical momentum and prioritizing the current gradient information. By adaptively modulating the momentum, ADM effectively dampens erratic gradient fluctuations while remaining responsive to significant directional shifts, yielding a more stable and reliable gradient $\widetilde{\mathbf{g}}_{l}$ for guiding the optimization towards data consistency.
The complete SPGD procedure is detailed in Algorithm 1. In our experiments, we extend the warm-up strategy throughout the entire reverse process, and set the total number of outer diffusion steps $T=100$ and warm-up steps $N=5$, aligning with common Number of Function Evaluations (NFEs) in existing literature.
4.3. Theoretical Analysis
Here, we provide theoretical support for our methodology. Our analysis highlights how the SPGD warm-up optimizes for data fidelity, distinguishing it from prior posterior sampling techniques.
Figure 3. High-level illustration of our proposed SPGD. (a) the standard reverse process (Eq. 8), and (b) our proposed SPGD (Algorithm 1), showing the warm-up phase with smoothed gradient to enhance restoration stability.
Proposition 4.1.
Let $L_{t}(\mathbf{x}_{t})=\frac{1}{2}\|\mathbf{y}-\mathcal{A}(\hat{\mathbf{x}}_{0}(\mathbf{x}_{t}))\|_{2}^{2}$ be the likelihood objective at diffusion timestep $t$. Assume the likelihood gradient $\mathbf{g}_{l}(\mathbf{x}_{t})=\nabla_{\mathbf{x}_{t}}L_{t}(\mathbf{x}_{t})$ is $L$-Lipschitz continuous with respect to $\mathbf{x}_{t}$ for some $L>0$. Consider the SPGD warm-up phase without momentum ($\beta=0$) using $N\geq 1$ steps with step size $\eta=\zeta/N$. If the step size satisfies $\eta<1/L$, then the state $\mathbf{x}_{t}^{(N)}$ obtained after $N$ warm-up steps guarantees a decrease in the likelihood objective:
(14) $L_{t}(\mathbf{x}_{t}^{(N)})\leq L_{t}(\mathbf{x}_{t}^{(0)})-\eta\sum_{j=0}^{N-1}\left(1-\frac{L\eta}{2}\right)\|\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})\|_{2}^{2}$
where $\mathbf{x}_{t}^{(0)}$ is the initial state at time $t$ before the warm-up.
Proof.
Please refer to Appendix B. ∎
Proposition 4.1 formally supports that the SPGD warm-up iterations progressively decrease the likelihood objective $L_{t}$, thereby enhancing the data consistency of the state $\mathbf{x}_{t}$ before denoising. In other words, $\mathbf{x}_{t}^{(N)}$ yields a better estimate $\hat{\mathbf{x}}_{0}$ with smaller likelihood objective value compared to using the initial state $\mathbf{x}_{t}^{(0)}$.
As conceptually illustrated in Figure 3, SPGD contrasts with standard posterior sampling methods like DPS (Chung et al., 2023). In DPS, the likelihood gradient $\mathbf{g}_{l}$ is computed based on the initial state $\mathbf{x}_{t}$ (equivalent to our $\mathbf{x}_{t}^{(0)}$), but the correction is applied only after the denoising step produces $\mathbf{x}_{t-1}^{\prime}$. This timing mismatch, where a gradient relevant to $\mathbf{x}_{t}$ corrects the transformed state $\mathbf{x}_{t-1}^{\prime}$, can potentially lead to instability and cause trajectories to deviate from the data manifold (Wu et al., 2023). On the contrary, SPGD performs iterative refinement directly on $\mathbf{x}_{t}^{(j)}$ before invoking the denoiser. By providing the denoising step with an input $\mathbf{x}_{t}^{(N)}$ that is already better aligned with the measurements $\mathbf{y}$, SPGD potentially facilitates a more stable update process and a more accurate subsequent generation.
Refer to captionInputDDRMDPSDiffPIRDPPSRED-DiffOursReference
Figure 4. Image restoration results with $\sigma_{y}=0.01$. Row 1: random inpainting, Row 2: SR ($\times$4), Row 3: Gaussian deblurring.
Table 1. Quantitative evaluation of image restoration tasks on FFHQ 256$\times$256 and ImageNet 256$\times$256 with $\sigma_{y}=0.01$.Best results are marked in bold, and second-best results are underlined.
Inpaint (random) Deblur (Gaussian) Deblur (motion) SR ($\times$ 4)
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
FFHQ
PnP-ADMM (Chan et al., 2016) 27.99 0.729 0.306 26.07 0.758 0.260 25.86 0.772 0.278 27.75 0.835 0.246
Score-SDE (Song et al., 2021b) 22.04 0.613 0.370 24.32 0.745 0.336 15.87 0.495 0.541 18.53 0.272 0.598
MCG (Chung et al., 2022) 27.74 0.825 0.147 24.28 0.737 0.292 19.78 0.540 0.476 18.10 0.261 0.606
DDRM (Kawar et al., 2022) 22.02 0.652 0.362 27.11 0.779 0.244 - - - 29.49 0.853 0.190
DPS (Chung et al., 2023) 26.11 0.802 0.180 26.51 0.782 0.181 25.58 0.752 0.212 27.06 0.803 0.187
DiffPIR (Zhu et al., 2023) 28.64 0.842 0.182 25.92 0.722 0.234 27.51 0.635 0.264 28.21 0.801 0.211
DPPS (Wu et al., 2024a) 25.47 0.769 0.191 26.92 0.793 0.173 26.25 0.771 0.194 27.52 0.820 0.153
RED-Diff (Mardani et al., 2024) 27.17 0.799 0.159 24.69 0.672 0.288 26.24 0.706 0.255 29.06 0.800 0.243
Ours 30.87 0.889 0.120 27.83 0.775 0.172 29.41 0.834 0.158 29.35 0.831 0.137
ImageNet
PnP-ADMM (Chan et al., 2016) 25.14 0.643 0.355 22.01 0.541 0.418 21.88 0.563 0.408 23.95 0.718 0.323
Score-SDE(Song et al., 2021b) 16.25 0.234 0.629 21.31 0.593 0.446 13.56 0.371 0.584 17.69 0.285 0.582
MCG (Chung et al., 2022) 23.21 0.647 0.294 12.31 0.340 0.578 18.32 0.425 0.502 17.08 0.264 0.591
DDRM (Kawar et al., 2022) 19.34 0.449 0.525 23.67 0.632 0.378 - - - 25.49 0.736 0.298
DPS (Chung et al., 2023) 25.09 0.773 0.217 19.49 0.485 0.399 18.68 0.454 0.432 23.47 0.665 0.310
DiffPIR (Zhu et al., 2023) 26.59 0.793 0.204 22.51 0.557 0.363 25.37 0.616 0.238 25.44 0.723 0.272
DPPS (Wu et al., 2024a) 24.55 0.755 0.193 21.75 0.587 0.352 20.09 0.530 0.392 24.36 0.718 0.253
RED-Diff (Mardani et al., 2024) 24.86 0.750 0.171 22.80 0.577 0.339 24.84 0.687 0.284 25.09 0.686 0.325
Ours 26.28 0.798 0.165 24.80 0.651 0.229 28.11 0.728 0.154 25.71 0.705 0.203
5. Experiments
5.1. Experimental Setup
Datasets & Metrics. We evaluate our proposed SPGD method and baseline approaches on 1,000 images from the FFHQ 256$\times$256 validation set (Karras et al., 2019) and 500 images from the ImageNet 256$\times$256 validation set (Russakovsky et al., 2015). The pre-trained diffusion models for FFHQ and ImageNet are obtained from (Chung et al., 2023) and (Dhariwal and Nichol, 2021), respectively, following established conventions.
Performance is assessed with quantitative metrics: Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) (Wang et al., 2004) for distortion evaluation, and Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018) for perceptual quality.
Baselines. We compare our method against several state-of-the-art diffusion-based image restoration techniques, including PnP-ADMM (Chan et al., 2016), Score-SDE (Song et al., 2022), MCG (Chung et al., 2022), DDRM (Kawar et al., 2022), DPS (Chung et al., 2023), DiffPIR (Zhu et al., 2023), DPPS (Wu et al., 2024a), and RED-Diff (Mardani et al., 2024). All experiments are conducted using the same pre-trained models and a fixed random seed to ensure fair comparisons.
Tasks. We consider four standard linear image restoration tasks: (1) Inpainting: Randomly mask 80% of the pixels across all RGB channels. (2) Gaussian Blur: Apply a Gaussian blur kernel of size $61\times 61$ with a standard deviation of $\sigma=3.0$. (3) Motion Blur: Use a motion blur kernel of size $61\times 61$ and intensity 0.5, following the procedure of Chung et al. (Chung et al., 2022). (4) Super-Resolution (SR) $\times$4: Perform 4×4$\times$4 × bicubic downsampling.
In all scenarios, Gaussian noise with a variance of $\sigma_{y}=0.01$ is added to the degraded measurements. For efficiency, all our ablation studies were conducted using 100 FFHQ test images.
Table 2. Ablation study on the effect of components.
Strategies Inpainting SR ($\times$ 4)
[…]
30.35 0.861 0.155 28.41 0.785 0.194
30.34 0.859 0.157 28.36 0.783 0.197
31.57 0.902 0.093 29.79 0.838 0.131
31.63 0.904 0.075 30.08 0.847 0.124
5.2. Quantitative Results
As summarized in Table 1, SPGD demonstrates consistently superior performance across the majority of tasks and metrics.
On the FFHQ dataset, our method achieves the best results for all metrics across all four degradation types. This indicates that our approach effectively minimizes distortion (high PSNR), preserves structural details (high SSIM), and generates results with superior perceptual similarity to the ground truth (low LPIPS) compared to the baseline methods on this dataset. The improvements are particularly significant in inpainting and motion deblurring, highlighting the robustness of our method in handling complex degradations.
On the more challenging ImageNet dataset, our method continues to outperform the baselines in most scenarios. It achieves the best SSIM and LPIPS scores for inpainting, Gaussian deblurring, and motion deblurring. For SR×$\times$4, our method obtains the best PSNR and LPIPS, and competitive SSIM scores, demonstrating its effectiveness on diverse image content. While DDRM achieves slightly higher SSIM for SR ($\times$4), the reliance on the singular value decomposition (SVD) of operator $\mathcal{A}$ renders their methodology inapplicable in many scenarios. Our method shows a marked advantage in LPIPS and frequently outperforms others in distortion metrics, demonstrating that our gradient management directly leads to the enhanced stability with superior restoration quality.
5.3. Qualitative Results
Visual comparisons, presented in Figure 4, corroborate the quantitative findings. Across different restoration tasks, SPGD consistently produces visually superior results compared to baseline methods. For instance, in random inpainting on FFHQ, SPGD generates significantly cleaner and more coherent reconstructions, successfully restoring intricate details like eye reflections and hair textures where competing methods often exhibit artifacts or blurriness. In super-resolution, SPGD renders sharper textures and finer structural details compared to the outputs of methods like RED-Diff, which can appear overly smoothed. Similarly, for Gaussian deblurring tasks, SPGD yields noticeably sharper edges and recovers finer details compared to DiffPIR or DPPS. These visual improvements are direct consequences of the stabilized reverse trajectory enabled by the proposed techniques.
Furthermore, we analyze the evolution of the predicted image ($\hat{\mathbf{x}}_{0}$) of the challenging task motion deblurring. As shown in Figure 5, the method DPS exhibits erratic behavior in intermediate stages (e.g., NFE = 150). Excessive or unstable likelihood corrections, driven by the fluctuating likelihood gradient $\mathbf{g}_{l}$, can introduce strong artifacts into $\hat{\mathbf{x}}_{0}$, potentially disrupting the global structure of the final result. In contrast, SPGD demonstrates a much smoother and more stable progression. The progressive warm-up and ADM smoothing enable gradual, controlled refinement guided by the likelihood, preserving key image features and more reliably leading to a perceptually convincing restoration, particularly in complex restoration scenarios. These results strongly support our analysis in Section 3.
Figure 5. Comparison of intermediate predictions $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})$ from DPS and our SPGD method on motion deblurring.
Figure 6. Evolution of PSNR and LPIPS of $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})$ during the reverse process for SPGD (T=100, N=5) and the baseline without gradient management (T=500).
5.4. Ablation Study
5.4.1. Impact of Design Components
SPGD integrates two core ideas: the progressive warm-up strategy and the ADM smoothing technique. To isolate their respective contributions, we evaluated our ablation study on four variants with 100 FFHQ test images: baseline [^2] The performance discrepancy between ’baseline’ and DPS is attributable to our improved experimental settings, as detailed in Appendix D.; SPGD with only the warm-up strategy; SPGD with only ADM smoothing; and the full SPGD model.
The results, presented in Table 2, reveal the following observations: (i) Implementing only the warm-up strategy yields significant improvements across all metrics compared to the baseline. This confirms that performing multiple, smaller likelihood updates before the denoising step effectively mitigates gradient conflicts and enhances reconstruction quality. (ii) Applying ADM smoothing directly fails to improve performance. We hypothesize this occurs because the momentum accumulates information based on likelihood gradients calculated at different states ($t-1$, $t$, etc.) where the underlying image estimate has significantly changed due to the intervening denoising steps. This makes the historical gradient information less relevant or potentially misleading for the current step’s likelihood correction. (iii) Combining both techniques results in the best performance, substantially outperforming either component alone. This demonstrates a crucial synergy: ADM effectively stabilizes the likelihood gradient within the warm-up phase, where likelihood gradients $\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})$ are computed at the same effective timestep $t$. These findings underscore that both components are essential and work synergistically, validating the design of SPGD.
5.4.2. Ablation Study on $N$
The number of inner warm-up steps, denoted as $N$, is a central hyperparameter in our proposed SPGD. To analyze its impact, we performed an ablation study varying $N$ while keeping the number of outer diffusion steps fixed at $T=100$. We evaluated $N\in\{2,5,10,20\}$ and included standard DPS with $T=500$ and $T=1000$ steps (corresponding to 500 and 1000 NFEs respectively) as references.
Results are presented in Table 3. We observe that increasing $N$ from 2 to 10 generally improves performance, particularly benefiting the perceptual LPIPS metric. Optimal performance appears around $N=5$ or $N=10$, depending on the specific metric. However, increasing $N$ further to 20 leads to a noticeable decline in quality, potentially due to overfitting within the warm-up phase. Considering the trade-off between performance and the computational overhead, we select $N=5$ as the default value for SPGD.
Notably, as shown in Table 3, with the same 500 NFEs, SPGD with $T=100,N=5$ significantly outperforms standard DPS using $T=500$ steps, and even surpasses DPS with $T=1000$ steps. Further evidence is provided in Figure 6, which compares the evolution of PSNR and LPIPS during the reverse process for SPGD (T=100, N=5) and DPS (T=500). SPGD consistently achieves superior metric values throughout the generation process for the same NFEs, indicating not only better final results but also faster convergence towards a high-quality solution.
Table 3. Ablation study on the impact of warm-up steps $N$.
[…]
$T$= 100 $m$=1 30.357 0.861 0.154
$T$= 100 $m$=2 31.655 0.896 0.109
$T$= 100 $m$=5 31.631 0.904 0.075
$T$= 100 $m$=10 30.923 0.893 0.070
$T$= 100 $m$=20 29.900 0.874 0.079
$T$=500 $m$=1 30.403 0.863 0.096
$T$=1000 $m$=1 31.608 0.902 0.080
5.4.3. Influence of $\beta$
To investigate the influence of the parameter $\beta$, we conduct an ablation study on Gaussian deblurring and SR $\times$4 tasks. Quantitative results, summarized in Table 4, reveal that setting $\beta$ within the range of 0.9 to 0.95 leads to optimal performance.
For the Gaussian Deblurring task, the configuration $\beta=0.9$ yields the highest PSNR (28.32), whereas $\beta=0.95$ achieves the best SSIM (0.788) and LPIPS (0.159) scores. In the case of SR $\times$4, $\beta=0.95$ consistently outperforms other values across all evaluation metrics. Notably, performance degrades as $\beta$ approaches 1.0, with the particularly decline in the deblurring scenario. Based on these observations, we adopt $\beta=0.95$ as the default setting in our method.
Table 4. Ablation study on the impact of parameter $\beta$.
Deblur (Gaussian) SR ($\times$4)
[…]
$\beta$=0.8 28.26 0.781 0.166 30.02 0.843 0.130
$\beta$=0.9 28.32 0.787 0.161 30.06 0.845 0.127
$\beta$=0.95 28.28 0.788 0.159 30.08 0.847 0.124
$\beta$=0.99 26.93 0.756 0.189 29.99 0.846 0.124
$\beta$=1.0 25.15 0.669 0.249 29.92 0.846 0.126
6. Conclusion
In this paper, we examined the reverse update from a gradient decomposition perspective, identified and addressed previously under-addressed gradient conflicts and fluctuations that hinder diffusion-based image restoration methods. Our proposed SPGD, incorporating progressive warm-up and adaptive momentum smoothing, directly mitigates these instabilities, leading to a more stable reverse process. Extensive experiments demonstrated SPGD’s superiority over state-of-the-art methods across various tasks, achieving significant quantitative and qualitative improvements, often with enhanced efficiency.
SPGD provides a practical improvement in image restoration and offers valuable insights into stabilizing guided diffusion processes. Future work could broaden the application of SPGD to diverse conditional diffusion frameworks, including text-to-image synthesis, image-to-image translation, and video generation.
Acknowledgements.
This work was supported by the National Major Scientific Instruments and Equipments Development Project of National Natural Science Foundation of China under Grant 62427820, the Fundamental Research Funds for the Central Universities under Grant 1082204112364, and the Science Fund for Creative Research Groups of Sichuan Province Natural Science Foundation under Grant 2024NSFTD0035.
Appendix A Derivation for Reverse Update Decomposition
In this part, we formally derive the DPS algorithm with DDIM sampling to our gradient perspective decomposition Equation 9.
The standard DDIM sampling with $\sigma_{t}=0$ is given by:
(15) $\mathbf{x}_{t-1}^{\prime}=\sqrt{\bar{\alpha}_{t-1}}\left(\frac{\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)+\sqrt{1-\bar{\alpha}_{t-1}}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t),$
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ is the cumulative product of the noise schedule, and $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ denotes the estimated noise obtained from the learned network.
Rewriting the scaling term,
$\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}=\frac{1}{\sqrt{\alpha_{t}}},$
we express $\mathbf{x}_{t-1}^{\prime}$ as:
(16) $\mathbf{x}_{t-1}^{\prime}$ $=\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}\mathbf{x}_{t}-\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)+\sqrt{1-\bar{\alpha}_{t-1}}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$
$=\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}\mathbf{x}_{t}-\left(\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}\sqrt{1-\bar{\alpha}_{t}}-\sqrt{1-\bar{\alpha}_{t-1}}\right)\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$
$=\frac{1}{\sqrt{\alpha_{t}}}\mathbf{x}_{t}-\left(\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\alpha_{t}}}-\sqrt{1-\bar{\alpha}_{t-1}}\right)\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t).$
Here, the second term represents the denoising gradient, which we denote as:
(17) $\mathbf{g}_{d}(\mathbf{x}_{t})=\left(\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\alpha_{t}}}-\sqrt{1-\bar{\alpha}_{t-1}}\right)\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t).$
For conditional generation based on measurement $\mathbf{y}$, DPS introduce an additional likelihood update:
(18) $\mathbf{x}_{t-1}=\mathbf{x}_{t-1}^{\prime}-\zeta\nabla_{\mathbf{x}_{t}}\|\mathbf{y}-\mathcal{A}(\hat{\mathbf{x}}_{0}(\mathbf{x}_{t}))\|_{2}^{2},$
where $\mathbf{y}$ represents the observed data, $\mathcal{A}$ is the observation operator, and $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})$ is the estimated clean image given by:
(19) $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})=\frac{\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}.$
The likelihood gradient term is thus defined as:
(20) $\mathbf{g}_{l}(\mathbf{x}_{t})=\nabla_{\mathbf{x}_{t}}\|\mathbf{y}-\mathcal{A}(\hat{\mathbf{x}}_{0}(\mathbf{x}_{t}))\|_{2}^{2}.$
Finally, the complete algorithm in an gradient perspective can be expressed as:
(21) $\mathbf{x}_{t-1}=$ $\underbrace{\frac{1}{\sqrt{\alpha_{t}}}\mathbf{x}_{t}}_{\text{fixed scaling}}-\underbrace{\left(\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{{\alpha}_{t}}}-\sqrt{1-\bar{\alpha}_{t-1}}\right)\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)}_{\text{denoising gradient}~{}\mathbf{g}_{d}(\mathbf{x}_{t})}$
$-{\zeta}\underbrace{\nabla_{\mathbf{x}_{t}}\|\mathbf{y}-\mathcal{A}(\hat{\mathbf{x}}_{0}(\mathbf{x}_{t}))\|_{2}^{2}}_{\text{likelihood gradient}~{}\mathbf{g}_{l}(\mathbf{x}_{t})}$
$=$ $\underbrace{\frac{1}{\sqrt{\alpha_{t}}}\mathbf{x}_{t}}_{\text{fixed scaling}}-\underbrace{\mathbf{g}_{d}(\mathbf{x}_{t})}_{\text{denoising gradient}}-\zeta\underbrace{\mathbf{g}_{l}(\mathbf{x}_{t})}_{\text{likelihood gradient}},$
which matches the Equation 9 in the main paper.
Appendix B Proofs
B.1. Descent Lemma for L-Smooth Functions
To support the proof of Proposition 4.1, we first state and prove the following lemma concerning functions with $L$-smooth gradients.
Lemma B.1.
Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be a continuously differentiable function whose gradient $\nabla f$ is $L$-Lipschitz continuous for some $L>0$, i.e., $\|\nabla f(\mathbf{w}^{\prime})-\nabla f(\mathbf{w})\|_{2}\leq L\|\mathbf{w}^{\prime}-\mathbf{w}\|_{2}$ for all $\mathbf{w},\mathbf{w}^{\prime}\in\mathbb{R}^{d}$. Then, for any $\mathbf{w},\mathbf{w}^{\prime}\in\mathbb{R}^{d}$:
(22) $f(\mathbf{w}^{\prime})\leq f(\mathbf{w})+\langle\nabla f(\mathbf{w}),\mathbf{w}^{\prime}-\mathbf{w}\rangle+\frac{L}{2}\|\mathbf{w}^{\prime}-\mathbf{w}\|_{2}^{2}$
Proof.
By the Fundamental Theorem of Calculus for vector functions, we have:
[…]
$f(\mathbf{w}^{\prime})-f(\mathbf{w})=$ $\langle\nabla f(\mathbf{w}),\mathbf{w}^{\prime}-\mathbf{w}\rangle+\int_{0}^{1}\langle\nabla f(\mathbf{w}+t(\mathbf{w}^{\prime}-\mathbf{w}))$
(24) $-\nabla f(\mathbf{w}),\mathbf{w}^{\prime}-\mathbf{w}\rangle\,dt$
Applying the Cauchy-Schwarz inequality to the integrand:
$\langle\nabla f(\mathbf{w}+t(\mathbf{w}^{\prime}-\mathbf{w}))-\nabla f(\mathbf{w}),\mathbf{w}^{\prime}-\mathbf{w}\rangle$
(25) $\leq$ $\|\nabla f(\mathbf{w}+t(\mathbf{w}^{\prime}-\mathbf{w}))-\nabla f(\mathbf{w})\|_{2}\|\mathbf{w}^{\prime}-\mathbf{w}\|_{2}$
Using the $L$-Lipschitz property of the gradient $\nabla f$:
$\|\nabla f(\mathbf{w}+t(\mathbf{w}^{\prime}-\mathbf{w}))-\nabla f(\mathbf{w})\|_{2}$
$\leq$ $L\|(\mathbf{w}+t(\mathbf{w}^{\prime}-\mathbf{w}))-\mathbf{w}\|_{2}$
$=$ $L\|t(\mathbf{w}^{\prime}-\mathbf{w})\|_{2}$
(26) $=$ $Lt\|\mathbf{w}^{\prime}-\mathbf{w}\|_{2}$
Substituting this back into the integral bound:
$f(\mathbf{w}^{\prime})-f(\mathbf{w})$
[…]
(28) $f(\mathbf{w}^{\prime})-f(\mathbf{w})\leq\langle\nabla f(\mathbf{w}),\mathbf{w}^{\prime}-\mathbf{w}\rangle+\frac{L}{2}\|\mathbf{w}^{\prime}-\mathbf{w}\|_{2}^{2},$
which gives the desired inequality. ∎
B.2. Proof of Proposition 4.1
Proof of Proposition 4.1.
We apply the Lemma B.1 to each iteration $j$ of the SPGD warm-up loop (Line 9 in Algorithm 1, with $\beta=0$). The update is $\mathbf{x}_{t}^{(j+1)}=\mathbf{x}_{t}^{(j)}-\eta\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})$. Let $f=L_{t}$, $\mathbf{w}=\mathbf{x}_{t}^{(j)}$, and $\mathbf{w}^{\prime}=\mathbf{x}_{t}^{(j+1)}$. The displacement is $\mathbf{w}^{\prime}-\mathbf{w}=-\eta\mathbf{g}_{l}(\mathbf{w})$. Substituting into Lemma B.1, we get:
$L_{t}(\mathbf{x}_{t}^{(j+1)})$ $\leq L_{t}(\mathbf{x}_{t}^{(j)})+\langle\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)}),-\eta\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})\rangle+\frac{L}{2}\|-\eta\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})\|_{2}^{2}$
$=L_{t}(\mathbf{x}_{t}^{(j)})-\eta\langle\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)}),\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})\rangle+\frac{L}{2}\eta^{2}\|\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})\|_{2}^{2}$
$=L_{t}(\mathbf{x}_{t}^{(j)})-\eta\|\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})\|_{2}^{2}+\frac{L\eta^{2}}{2}\|\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})\|_{2}^{2}$
(29) $=L_{t}(\mathbf{x}_{t}^{(j)})-\eta\left(1-\frac{L\eta}{2}\right)\|\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})\|_{2}^{2}.$
The condition $\eta<1/L$ ensures that $(1-L\eta/2)>1/2>0$. Thus, each warm-up step guarantees a non-increasing likelihood objective, $L_{t}(\mathbf{x}_{t}^{(j+1)})\leq L_{t}(\mathbf{x}_{t}^{(j)})$, with strict decrease if $\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})\neq\mathbf{0}$.
Summing this per-step inequality over $j=0,\dots,N-1$ yields:
(30) $\sum_{j=0}^{N-1}[L_{t}(\mathbf{x}_{t}^{(j+1)})-L_{t}(\mathbf{x}_{t}^{(j)})]\leq\sum_{j=0}^{N-1}\left[-\eta\left(1-\frac{L\eta}{2}\right)\|\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})\|_{2}^{2}\right]$
The right-hand side can be simplified by moving the constant factors outside the summation:
$\sum_{j=0}^{N-1}[L_{t}(\mathbf{x}_{t}^{(j+1)})-L_{t}(\mathbf{x}_{t}^{(j)})]\leq-\eta\left(1-\frac{L\eta}{2}\right)\sum_{j=0}^{N-1}\|\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})\|_{2}^{2}$
The left-hand side of Eq. (30) is a telescoping sum:
$\sum_{j=0}^{N-1}[L_{t}(\mathbf{x}_{t}^{(j+1)})-L_{t}(\mathbf{x}_{t}^{(j)})]$
$=$ $[L_{t}(\mathbf{x}_{t}^{(1)})-L_{t}(\mathbf{x}_{t}^{(0)})]+[L_{t}(\mathbf{x}_{t}^{(2)})-L_{t}(\mathbf{x}_{t}^{(1)})]+\dots$
$\dots+[L_{t}(\mathbf{x}_{t}^{(N)})-L_{t}(\mathbf{x}_{t}^{(N-1)})]$
$=$ $L_{t}(\mathbf{x}_{t}^{(N)})-L_{t}(\mathbf{x}_{t}^{(0)})$
Substituting this back, we have:
$L_{t}(\mathbf{x}_{t}^{(N)})-L_{t}(\mathbf{x}_{t}^{(0)})\leq-\eta\left(1-\frac{L\eta}{2}\right)\sum_{j=0}^{N-1}\|\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})\|_{2}^{2}$
Rearranging this inequality to isolate $L_{t}(\mathbf{x}_{t}^{(N)})$ yields the final result stated in the proposition:
$L_{t}(\mathbf{x}_{t}^{(N)})\leq L_{t}(\mathbf{x}_{t}^{(0)})-\eta\sum_{j=0}^{N-1}\left(1-\frac{L\eta}{2}\right)\|\mathbf{g}_{l}(\mathbf{x}_{t}^{(j)})\|_{2}^{2}$
This completes the proof. ∎
Appendix C More Discussions
C.1. Novel Aspects of SPGD
While prior works have explored momentum-based techniques (He et al., 2024; Chung et al., 2024) and multi-step likelihood updates (Xu et al., 2025; Ye et al., 2024b; Huang et al., 2024), SPGD’s novelty lies in three key areas:
  1. (1)
    Analysis-Driven Motivation: Our core contribution is the novel gradient-centric analysis identifying the dual problems of gradient conflict and gradient fluctuation. SPGD’s components are not an ad-hoc combination but are directly motivated by this analysis to solve these specific, identified issues, providing a principled foundation.
  2. (2)
    Technical Novelty of ADM: Unlike prior works that use standard, constant momentum, our ADM is technically more advanced. It dynamically adjusts the momentum strength based on directional consistency (Eq. (13)), making it more robust than a fixed momentum coefficient, as validated in our ablation studies.
  3. (3)
    Synergistic Design: Our warm-up phase is specifically designed to mitigate gradient conflict before the denoising step. This creates a stable context that enables ADM to effectively dampen fluctuations. The synergistic combination of a conflict-mitigating warm-up with a fluctuation-dampening ADM is a unique design absent in prior art.
Refer to captionbaselinew/ ADMw/ warm-upSPGDreference
Figure 7. Visual results of ablation study on design components. The complete SPGD shows clearer hair textures.
C.2. Relation to Pixel-as-Parameter (PxP)
Our SPGD method shares conceptual ground with Pixel-as-Parameter (PxP) (Dinh et al., 2023) in viewing the reverse process through a gradient lens and acknowledging potential gradient conflicts. However, our solutions are fundamentally different.
PxP’s approach is to explicitly resolve conflicts by projecting gradients to enforce orthogonality before the update. In contrast, SPGD aims to mitigate conflicts by using a progressive warm-up phase. This phase allows the state $\mathbf{x}_{t}$ to gradually adapt to the likelihood guidance before the denoising step is applied. We adopt this strategy because we found that explicit gradient projection can lead to information loss and degrade performance.
Furthermore, PxP computes all guidance terms from the initial $\mathbf{x}_{t}$ and applies them in a single, combined update. SPGD’s architecture is different, introducing an inner loop that iteratively refines $\mathbf{x}_{t}$ with the likelihood gradient prior to denoising. In essence, SPGD employs sequential refinement for conflict mitigation, whereas PxP uses simultaneous projection for conflict resolution.
C.3. ADM Stability
While a full theoretical convergence proof for SPGD with ADM is a significant undertaking beyond the scope of this paper, we can provide further intuition on its stability. The ADM update rule, $\widetilde{\mathbf{g}}_{l}^{(j)}=\alpha_{j}\beta\widetilde{\mathbf{g}}_{l}^{(j-1)}+(1-\alpha_{j}\beta)\mathbf{g}_{l}^{(j)}$, can be interpreted as a standard momentum update with a dynamically adjusted momentum coefficient $\beta^{\prime}=\alpha_{j}\beta$.
The stability of standard momentum methods is well-understood and often relies on the step size and momentum coefficient being within certain bounds related to the properties of the objective function (e.g., its smoothness $L$ and strong convexity $\mu$). Our adaptive weight $\alpha_{j}\in[0,1]$ ensures that the effective momentum $\beta^{\prime}$ is always less than or equal to the base coefficient $\beta$ ($\beta^{\prime}\leq\beta$). When the gradient direction changes abruptly, $\alpha_{j}$ decreases, reducing the effective momentum $\beta^{\prime}$. This reduction pushes the update behavior closer to that of standard gradient descent, a method with well-known and robust convergence guarantees under standard assumptions (e.g., sufficiently small step size).
In essence, ADM acts as a self-stabilizing mechanism. It leverages the acceleration benefits of momentum when the optimization landscape is smooth and consistent, but reduces its reliance on momentum precisely when the gradient signal becomes noisy or changes direction sharply. This connection to the more conservative gradient descent method provides strong theoretical intuition for why ADM contributes to the overall stability of the SPGD framework.
Appendix D Implementation Details
D.1. Comparison Methods
To ensure a fair comparison, all pixel-based diffusion methods (Score-SDE, MCG, DDRM, DPS, DiffPIR, DPPS, and RED-Diff) share the same pre-trained checkpoint. Below, we describe our implementation details for each method.
PnP-ADMM: We use the pre-trained DnCNN model from (Zhang et al., 2017). The regularization parameter is set to $\tau=0.2$, and the algorithm is run for 10 iterations on all tasks.
Score-SDE: In Score-SDE, data consistency is enforced via a projection step following unconditional diffusion denoising at every iteration. The projection configuration follows the recommendations in (Chung et al., 2022).
MCG: Our implementation of MCG is based on the source code provided in (Chung et al., 2022). We replicate the evaluation settings reported in the paper.
DPS: Experimental outcomes for DPS are obtained from the implementation in (Chung et al., 2023) using 1,000 DDPM sampling steps. All hyper-parameters are set as suggested in D.1 section of the corresponding paper.
DDRM: For DDRM, we utilize 20 NFEs DDIM sampling (Song et al., 2021a) with parameters $\eta=0.85$ and $\eta_{B}=1.0$, as recommended in the literature.
DiffPIR: We adopt the original code of DiffPIR from (Zhu et al., 2023). We adjust the diffusion steps to 20 for implementation efficiency. Hyper-parameters are selected to reflect the settings for the noiseless case.
DPPS: We follow the default configuration as provided by the open-sourced code in (Wu et al., 2024a).
RED-Diff: The implementation for RED-Diff is adopted from (Mardani et al., 2024). We use the Adam optimizer over 1,000 steps, setting $\lambda=0.25$ and $lr=0.25$, consistent with the experimental settings in the paper.
D.2. Our SPGD:
For our SPGD implementation detailed in Algorithm 1, we follow the time step discretization and noise schedule from EDM (Karras et al., 2022). We use $T=100$ outer diffusion steps. Each outer step includes $N=5$ inner warm-up iterations, corresponding to 500 NFEs in total. The base momentum coefficient for ADM is set to $\beta=0.95$.
The learning rate $\zeta$ is tuned separately for each task. The specific values are provided in Table 5.
Table 5. Learning rates $\zeta$ for each task.
Dataset FFHQ ImageNet
Inpainting 2.5 2.5
Deblurring (Gaussian) 1.5 1
Deblurring (Motion) 1 1
Super-Resolution (×$\times 4$) 8 7
D.3. Computational Cost
Table 6. Computational Cost ($T=100$) under different $N$.
Metric $N$=1 $N$=2 $N$=5 $N$=10 $N$=20
FLOPS (T) ↓ 19.38 38.75 96.89 193.77 387.54
Time (s) ↓ 8.69 15.11 33.83 65.19 129.14
To provide a clearer understanding of the computational overhead, we report both the inference time and the forward-pass cost (in FLOPs) of the network $\bm{\epsilon}_{\bm{\theta}}$ for each image on RTX 4090 GPU, with $T=100$ and varying $N$ on FFHQ inpainting.
Appendix E More Visual Results
In this section, we provide supplementary experimental results to further substantiate the effectiveness of our proposed method. Specifically, Figure 7 presents high-resolution visualizations for the ablation study on the design components. Figure 8 to Figure 11 further illustrate the visual advantages of our proposed approach. Figure 12 to Figure 15 provide additional examples of angular relationships of gradients, both before and after ADM smoothing.
Figure 8. Qualitative visual comparison of random inpainting on FFHQ (top rows) and ImageNet (bottom rows).
Figure 9. Qualitative visual comparison of SR $\times$4 on FFHQ (top rows) and ImageNet (bottom rows).
Figure 10. Qualitative visual comparison of Gaussian deblurring on FFHQ (top rows) and ImageNet (bottom rows).
Figure 11. Qualitative visual comparison of motion deblurring on FFHQ (top rows) and ImageNet (bottom rows).
(a) Sample 1
(b) Sample 2
(c) Sample 3
(d) Sample 4
(e) Sample 5
(f) Sample 6
Figure 12. Gradient angles dynamics before (left) and after ADM smoothing (right) on random inpainting.
(a) Sample 1
(b) Sample 2
(c) Sample 3
(d) Sample 4
(e) Sample 5
(f) Sample 6
Figure 13. Gradient angles dynamics before (left) and after ADM smoothing (right) on Gaussian deblurring.
(a) Sample 1
(b) Sample 2
(c) Sample 3
(d) Sample 4
(e) Sample 5
(f) Sample 6
Figure 14. Gradient angles dynamics before (left) and after ADM smoothing (right) on motion deblurring.
(a) Sample 1
(b) Sample 2
(c) Sample 3
(d) Sample 4
(e) Sample 5
(f) Sample 6
Figure 15. Gradient angles dynamics before (left) and after ADM smoothing (right) on SR $\times$4.